1. Calculate the approximate [OH-] and [NH4+] in a 0.40 M ammonia solution, NH3(aq).

NH3(aq) + H2O(l) ↔ OH-(aq) + NH4+(aq). Kb = 1.75 x 10-5M.

2. Calculate the pH of 0.178 M ammonia.

NH3(aq) + H2O(l) ↔ OH-(aq) + NH4+(aq) Kb = 1.75 x 10-5
Answer #1

1)

NH3 + H2O ----> OH- + NH4 + , Kb = 1.75*10^-5

0.40 0 0 (at t=0)

0.40-x x x (at equilibrium)

Kb = x*x / 0.40-x

1.75*10^-5 = x^2 / 0.40-x

0.000007 - 0.0000175x = x^2

x^2 + 0.0000175x - 0.000007 = 0

Solving above quadratic equation,

x = 0.00264

[OH-] = [NH4+] = x = 0.00264 M ..... Answer

2) Similarly as above,

Kb = x^2 / 0.178-x

0.0000175 * (0.178-x) = x^2

0.000003115 - 0.0000175x = x^2

x^2 + 0.0000175x - 0.000003115 = 0

x = 0.00176

[OH-] = [NH4 +] = x = 0.00176 M .... Answer

Let me know in case of any doubts.